OpenStudy (anonymous):

Solve. x2 = 6x – 6 a.) x = 3 ± sqrt 3 b.) x = 6 ± sqrt 3 c.) x = 3 ± sqrt 30 d.) x = 3 ± sqrt 6

OpenStudy (anonymous):

Here is the steps to solve it.

OpenStudy (sburchette):

First manipulate the equation until all the terms are on one side. \[x^2-6x+6=0\] Then use the quadratic formula to obtain the solutions because it cannot be factoring using trinomial techniques. \[(-b \pm \sqrt{b^2-4ac})/2a\] Based on this quadratic equation give, a=1, b=-6, and c=6 Substituting we get \[(6\pm \sqrt{(-6)^2-4(1)(6)})/2(1)\]This simplifies to\[(6\pm \sqrt{12})/2\]Which further simplifies to \[3\pm \sqrt{3}\]

OpenStudy (anonymous):

Thank you! Much appreciation.

OpenStudy (sburchette):

You could have used completing the square, but using the quadratic formula is essentially completing the square.